-7.0 με 2.0 με 0.4 m 4.0 με 0.9 m

What is the magnitude of the net electric field at the fourth corner of the rectangle (the lower right)?

What is the angle that the electric field at that corner of the rectangle makes with the horizontal?

If we placed a small bead with charge 1.5 μC at that corner, what is the magnitude of the net electric force it would experience?

If we instead placed a small bead with charge -3.0μC at that corner, what is the magnitude of the net electric force it would experience?
Answer #1

Eure 9x10 Yuxi) fo. - 44444.44 Na - 9x10? (2x ) - Kar 24 = - coul - 112500 Nk i - no 4 to 8 20.97 nant & tim! ( 0 ) = 23.56

2 tail/-86124.52 - 14907.35 - 60.120 (80.17 below the - nanis or 180 +80.17 = 260017 above the tra cole 6 = Eg = 67405.86 (1-